OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -one thirdx + 7. y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

Well, from the choices we can guess that the slope of that line is going to be 3> We can verify by taking -1/3, flipping it and changing the sign. So m = 3. At this point, we should use the point slope formula. \[y-y _{0}=m(x-x _{0})\]

OpenStudy (anonymous):

y-3=3(x+1) y=3x+6

OpenStudy (anonymous):

So: y - 3 = 3 (x- (-1) y - 3 = 3x + 3 y = 3x + 6